# Modelling the Spread of New Information on Social Networks

### v3

@@ Abstract
There has been considerable interest in modelling the spread of information on social networks using machine learning models. Here, we consider the problem of predicting the spread of new information, i.e. when a user propagates information about a topic previously unseen by the user. In existing work, information and users are randomly assigned to a test or training set, ensuring that both sets are drawn from the same distribution. In the spread of new information, the problem becomes an out-of-distribution generalisation classification task.
Our experimental results reveal that while existing algorithms, which predominantly use features derived from the content of messages, perform well when the training and test distributions are the same, these algorithms perform much worse when the test set is out-of-distribution, i.e. when the topic (hashtag) of the testing data is absent from the training data. We then show that if the message features are supplemented or replaced with features derived from users’ profile and past behaviour, the out-of-distribution prediction is greatly improved, with the F1 score increasing from 0.117 to 0.705. Our experimental results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from users’ profile and past behaviour, and is largely content-agnostic.
Keywords Online Social Media $\cdot$ Social Content Sharing $\cdot$ Repost/Retweet $\cdot$ Out-of-Distribution Generalisation $\cdot$ Prediction $\cdot$ Machine Learning $\cdot$ X (Twitter)
@@ 1 Introduction
Understanding and predicting user reposting behaviour on social media platforms is an area of interest across disciplines such as computer science, social science, political science, and marketing. This interest stems from the influence that individual behaviours have on information diffusion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], including how information is spread during marketing and political campaigns [4, 5, 6, 7, 5, 10], and how misinformation and rumours spread [9, 8, 7, 10, 12].
@@
The prediction of reposting has been formulated as a supervised classification task. The objective is to train a binary classifier to estimate the likelihood of a recipient reposting a message from a sender. In this setting, the dataset is typically divided into training and test sets through random assignment, ensuring that both sets are drawn from the same underlying data distribution for fair and stable evaluation.
@@
On social media platforms, new topics appear regularly, e.g. those related to breaking news. Thus, in addition to the typical in-distribution classification, out-of-distribution prediction is also needed, where a model is trained on existing data related to previous topics (defined by hashtags) and is then used to predict a user’s reaction to a message related to a new, previously unseen, topic. Such a model can be useful for many applications, and it is more challenging, as it involves handling a distributional shift between user groups of known and unknown topics.
@@
In this paper, we first review previous work on predicting repostings on X (formerly Twitter) [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 3, 37, 38, 39, 2, 40, 41, 42, 1, 43, 44, 45, 46, 47, 48]. Existing algorithms predominantly focused on the textual data, including content of reposted messages and user historical posts. We argue that features of the sender and the recipient are also relevant to the reposting of a message. We study two types of data: (i) the message data, which contain the textual content of messages under consideration for reposting, (ii) the user data, which contain the sender and recipient’s profiles, their following lists, and their historical posts encoding their past activities, interests, and interactions with other users. We identified a total of 303 features from these data, including 78 message-related features and 225 user-related features. We consider three classification models: a decision tree model, a neural network model, and the BERT model [49], which is a widely studied Large Language Model (LLM). Using data collected on X, we conduct experiments for both in-distribution and out-of-distribution predictions.
@@
Our results show that while the prediction models, using only message-related features, perform well for in-distribution prediction, they perform much worse for out-of-distribution prediction. When the message-related features are supplemented or replaced with user-related features, the out-of-distribution performance greatly improves with the F1 score increasing from 0.117 to 0.705.
@@
Our work highlights the importance of user-related features for reposting prediction—features that are currently underutilized. Interestingly, our experimental results suggest that a significant component of reposting behaviour for previously unseen topics is largely content-agnostic. This work provides novel insights into the mechanisms underlying information spreading on social media.
@@ 2 Background
2.1 Reposting on Social Media Platforms
X (formerly Twitter) is one of the most popular social media platforms. It allows users to ‘post’ short messages (also called tweets), follow other users to curate their content feed, and interact with other users through actions, such as ‘like’ and ‘repost’. When a user posts a message, the message is shown to the user’s followers as well as other users who might be interested as determined by X’s recommendation algorithm.
A user can repostrepost (or ‘retweet’) a message. The act of reposting is often considered as an endorsement of a message and/or its creator. Reposting plays an important role in the diffusion of information on social media. On X, a user can also ‘quote’ or ‘reply’ to a message. Similarly to a repost, these actions will add the message to the timelines of the user’s followers and, therefore, will have an equivalent function as the repost in terms of information diffusion. In our X data, reposts account for 80% of the messages, and quotes and replies account for 2%. For simplicity, in this study we merge quotes and replies into reposts. A user can also ‘like’ a message to signal their endorsement, but this action does not add the message to the timelines of the user’s followers and has limited impact on information diffusion. Thus, the action of ‘like’ is not considered in this study.
Facebook and Sina Weibo also have similar repost functions (referred to as ‘share’ on Facebook). Despite the availability of datasets from these platforms, most research in literature has primarily focused on X due to its widespread use and greater accessibility until recently. Although Instagram has a ‘share’ function and TikTok offers a ‘repost’ feature, these differ from the ‘repost’ action discussed in this study, as they typically involve private forwarding of messages or posts to specific users or group chats.
@@ 2.2 Reposting Prediction
In this study, reposting prediction is defined as to predict whether a recipient will repost a message received from a sender. A reposting action involves three elements: (1) a message (M); (2) a sender ($U_{S}$) of the message who creates the message; and (3) a recipient ($U_{R}$) of the message via following or recommendation.
@@
$f:\{M,U_{S},U_{R}\}\rightarrow y\in\{0,1\}\,$, (1)
where $y=1$ means the recipient will repost the message received from the sender, and $y=0$ otherwise.
Table 1: Message and User Data and Features.
Data Feature Type Feature Sub-type Example feature Note
(# of features) (# of features)
Message Data Message-Related Features
Data(M) [0.1cm] Text content [0.1cm] of message M (78) Topic (39) M_TopicM5 Likelihood of a message related to topic ‘family’.
Language (10) M_Grammar1 Grammatical correctness of a message.
Readability (11) M_Readability1 A measure of difficulty in understanding a message.
Sentiment (5) M_Sentiment1 Negative sentiment score of a message.
Emotion (8) M_Emotion1 Probability of a message expressing emotion ‘anger’.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of a message.
Hashtag (1) M_Hashtag The presence of a specific hashtag in a message.
User Data User-Related Features for the Sender $U_{S}$ and the Recipient $U_{R}$
Data(U-P) [0.1cm] Profile and [0.1cm] friend list of [0.1cm] $U_{S}$ and $U_{R}$ U-P (30) Profile of $U_{S}$ (12) U-P_S_FollowerNum The number of followers of the Sender $U_{S}$.
Profile of $U_{R}$ (12) U-P_R_FollowerNum The number of followers of the Recipient $U_{R}$.
Network of $U_{S}$ (3) U-P_S_LeaderRank A measure of the Sender $U_{S}$’s influence.
Network of $U_{R}$ (3) U-P_R_LeaderRank A measure of the Recipient $U_{R}$’s influence.
Data(U-HA) [0.1cm] Public metrics [0.1cm] of historical [0.08cm] messages of [0.05cm] $U_{S}$ and $U_{R}$ U-HA (38) Popularity of $U_{S}$ (4) U-HA_S_LikedRate Avg. # of ‘Likes’ received by latest 50 messages of $U_{S}$.
Popularity of $U_{R}$(4) U-HA_R_LikedRate Avg. # of ‘Likes’ received by latest 50 messages of $U_{R}$.
Activity of $U_{S}$ (7) U-HA_S_InteractivePer Percentage of reposts in $U_{S}$’s latest 50 messages.
Activity of $U_{R}$(7) U-HA_R_InteractivePer Percentage of reposts in $U_{R}$’s latest 50 messages.
Interaction (16) U-HA_RS_Mention Number of $U_{R}$’s latest 50 messages that mentioned $U_{S}$.
Data(U-HM) U-HM (157) M features of $U_{S}$ (78) U-HM_S_Grammar1 (Each M feature is averaged over $U_{S}$’s latest 50 messages.)
Text content M features of $U_{R}$ (78) U-HM_R_Grammar1 (Each M feature is averaged over $U_{R}$’s latest 50 messages.)
of historical Topic Similarity (1) U-HM_SR_TopicSim Similarity between historical topics of 50 latest messages
messages of of $U_{S}$ and $U_{R}$.
$U_{S}$ and $U_{R}$
2.3 Existing Works
2.3.1 Related Works
Many works on the broader topic of information diffusion are related to the reposting action on social media. Some of them do not consider all the three elements defined in Equation (1), namely a message, a sender and a recipient. Some study messages only, some focus on messages and their senders only, and some consider messages and their recipients (i.e. reposters) only. For example, [26, 27] predicted the popularity of a message; [28, 29] predicted the incremental size of spreading tree for a given message over time; [30] predicted the time gap between reposts of a given message; [31] predicted the probability of a given message being reposted (by any user), and [32, 33, 34, 35, 36] predicted the next user who is going to repost a given message based on the sequence of previous reposters of the message and the social relations among these reposters.
2.3.2 Reposting Prediction
In this study, we focus on existing works on the reposting prediction task as defined in Equation (1) involving all the three elements.
These include a number of machine learning (ML) models, such as Support Vector Machines (SVM) [15, 16], Decision Trees  [15, 17, 13, 18], Logistic Regression  [19, 20], and factorisation models  [21, 22, 23, 24]. These models are based on features extracted from raw data for prediction. It is suggested in [15, 13] that the Decision Tree models often outperform other ML models.
@@ 2.3.2 Reposting Prediction
2.3.3 Limitations of Previous Reposting Prediction
Reliance on Message-Related Data
Previous models primarily relied on data of messages, including messages being reposted and the historial posts of users. For example, the ML models typically used (a relatively small set of) features mostly extracted from the textual content of message data. The NN models used the message data exclusively. The Bi-LSTM model even ignored the User ID associated with a message. It is known, however, that user-related data, including users profile, their following list, and their past behaviour, interest and interactions, can have a significant influence on a user’s status, actions and decisions [50, 51, 16, 15]. For example, [16] suggested that a recipient is more likely to repost a message from a sender if the recipient follows the sender, and the two users have shared interests and prior interactions.
Lack of Out-of-Distribution Evaluation
Existing works typically evaluated model performance by splitting the data into training and test sets using random sampling [15, 16, 17, 13, 18, 19, 21, 22, 23, 24, 14, 25], where both sets contain data from the same set of topics. This setting, often referred to as in-distribution evaluation [52, 53], assumes that the training and test data are drawn from the same underlying distribution. However, different topics may associate with user groups with distinct feature distributions. None of prior works explicitly evaluated generalisation across different topics in the out-of-distribution setting, that is to train a model using data of some topics and then test the model using data of other topics, which can evaluate a model’s capacity in terms of predicting reposts associated with new, previously unseen information.
3 Method
@@ 3.1 X Data
We collected data from X using the former Twitter API with the Academic Access. We selected the 14 most popular hashtags in the trending list between July 27th and August 14th in 2022 and retrieved all available messages containing these hashtags during the same period.
We obtained 111,401 messages, each containing one of the 14 trending hashtags. From these messages, we identified 44,014 reposts (positive instances) that satisfied the following conditions: (1) the original message being reposted was also in our data and (2) the recipient created the repost within 24 hours since the original message was created by the sender. According to the study of half-life of X messages [54], users are unlikely to see a message more than 24 hours after its creation. Any reposting after 24 hours are mostly influenced by external factors, such as TV or other social media platforms.
These messages were posted by 79,707 unique users around the world. For each user, we collected their profile, following list, and up to 50 latest historical messages. A total of more than 3.8 million historical messages were collected. All users are anonymised in our data.
3.2 Message Data and Features
A message can have text and multimedia content, such as pictures, audio and video. Same as in prior works, in this study, we only consider the text content of a message.
In this study, the message data, Data(M), contain the text content of the 111,401 messages in our X data. From the message data, we extract 78 message-related features, M, which are used in prior works for reposting prediction.
These features characterise a message from different perspectives. As shown in Table 1, we divide them into sub-types, including topic  [43, 41, 40, 13, 17, 24], language style [44, 21, 42, 2], readability [42], sentiment [45, 3, 38, 17], emotion [1, 3, 37, 17], hate speech  [46], and hashtag. These features are calculated using publicly available language models and tools (see Appendix). In addition, the hashtag feature is recorded as a categorical label [21, 17].
@@ 3.3 User Data and Features
In this study, the user data, Data(U), contains the following three types of data related to a user.
@@
  • Data(U-HA): the metadata (or public metrics) of the latest 50 historical messages of a user, including (1) the type of a message, i.e. post/repost/quote/reply, (2) the creation time of a message, (3) the number of reposts/quotes/replies/likes a message has received, and (4) any usernames mentioned in a message. These metadata encode useful information of a user’s previous behaviours, activity patterns, and its interactions with other users.
  • Data(U-HM): the textual content of the latest 50 historical messages of a user, which characterise a user’s historical interests and communication style.

### v4

# Modelling the Spread of New Information on X

@@ Abstract
There has been considerable interest in modelling the spread of information on X (formerly Twitter) using machine learning models. Here, we consider the problem of predicting the reposting of new information, i.e., when a user propagates information about a topic previously unseen by the user. In existing work, information and users are randomly assigned to a test or training set, ensuring that both sets are drawn from the same distribution. In the spread of new information, the problem becomes an out-of-distribution classification task.
Our experimental results reveal that while existing algorithms, which predominantly use features derived from the content of posts, perform well when the training and test distributions are the same, they perform much worse when the test set is out-of-distribution, i.e., when the topic of the testing data is absent from the training data. We then show that if the post features are supplemented or replaced with features derived from user profiles and past behaviours, the out-of-distribution prediction is greatly improved, with the F1 score increasing from 0.117 to 0.705. Our experimental results suggest that a significant component of reposting behaviour for previously unseen topics can be predicted from user profiles and past behaviours, and is largely content-agnostic.
Keywords Online Social Media $\cdot$ Social Content Sharing $\cdot$ Repost/Retweet $\cdot$ Out-of-Distribution Generalisation $\cdot$ Prediction $\cdot$ X (Twitter)
@@ 1 Introduction
Understanding and predicting user reposting behaviour on social networks is an area of interest across disciplines such as computer science, social science, political science, and marketing. This interest stems from the influence that individual behaviours have on information diffusion [40, 44, 39], including how information is spread during marketing [46] and political campaigns [47, 17, 13], and how misinformation and rumours spread [5, 8, 45].
@@
The prediction of reposting has been formulated as a supervised classification task. The objective is to train a binary classifier to estimate the likelihood of a recipient reposting a post from a sender. In this setting, the dataset is typically divided into training and test sets through random assignment, ensuring that both sets are drawn from the same underlying data distribution for fair and stable evaluation.
@@
On social media platforms, new topics appear regularly, e.g. those related to breaking news. Thus, in addition to the typical in-distribution classification, out-of-distribution prediction is also needed, where a model is trained on existing data related to previous topics (defined by hashtags) and is then used to predict a user’s reaction to a post related to a new, previously unseen, topic. Such a model can be useful for many applications, and it is more challenging, as it involves handling a distributional shift between user groups of known and unknown topics.
@@
In this paper, we first review previous work on predicting repostings on X (formerly Twitter). Existing algorithms predominantly focused on the textual data, i.e., the text content of posts. We argue that features of the sender and the recipient are also relevant to the reposting. We study two types of data: (i) the post data, which contain the textual content of posts under consideration for reposting, (ii) the user data, which contain the sender and recipient’s profiles, their following lists, and their historical posts encoding their past activities, interests, and interactions with other users. We identified a total of 303 features from these data, including 78 post-related features and 225 user-related features. We consider three classification models: a decision tree model, a neural network model, and the BERT model [14], which is a widely studied Large Language Model (LLM). Using data collected from X, we conduct experiments under both in-distribution and out-of-distribution settings.
@@
Our results show that while the prediction models, using only post-related features, perform well for in-distribution prediction, they perform much worse for out-of-distribution prediction. When the post-related features are supplemented or replaced with user-related features, the out-of-distribution performance greatly improves with the F1 score increasing from 0.117 to 0.705.
@@
Our work highlights the importance of user-related features for reposting prediction—features that are currently underutilized. Interestingly, our experimental results suggest that a significant component of reposting behaviour for previously unseen topics is largely content-agnostic. This work provides novel insights into the mechanisms underlying information spreading on social networks.
@@ 2 Background
X (formerly Twitter) is a popular online social network. It allows users to publish short posts (formerly known as tweets), follow other users to curate their content feed (so called ‘timeline’), and interact with other users through actions, such as ‘repost’ and ‘like’. When a user creates a post, the post will be added to the timelines of the user’s followers as well as those who might be interested as determined by X’s recommendation algorithm.
When a user receives a post, it can ‘repost’ (formerly known as retweet) the post, ‘reply’ to the post by adding their response, or ‘quote’ the post by adding their comment. In our X data, reposts account for 78% of all posts, and replies and quotes account for 2%. In this study, we merge replies and quotes into reposts, because these actions add the post to the timelines of their followers and therefore have the same effect on information diffusion. A user can also ‘like’ a post to signal their endorsement, but this action does not add the post to the timelines of the user’s followers. The reposting function on X is similar to the ‘share’ function on Facebook. The ‘share’ function on Instagram and the ‘repost’ function on TikTok, however, are different as they forward posts privately to specific users or groups.
Source Type(#) Sub-type(#) Examples Note
Post-Related Features
Data(M) Text content of post M (78) Topic (39) M_TopicM5 Likelihood of a post related to topic ‘family’.
Language (10) M_Grammar1 Grammatical correctness of a post.
Readability (11) M_Readability1 A measure of difficulty in understanding a post.
Sentiment (5) M_Sentiment1 Negative sentiment score of a post.
Emotion (8) M_Emotion1 Probability of a post expressing emotion ‘anger’.
Hate-Speech (4) M_Hate1 A measure of aggressiveness of a post.
Hashtag (1) M_Hashtag The presence of a specific hashtag in a post.
User-Related Features for Sender $U_{S}$ and Recipient $U_{R}$
Data(U-P) U-P (30) Profile of $U_{S}$ (12) U-P_S_FollowerNum The number of followers of the Sender $U_{S}$.
Profile and friend list of $U_{S}$ and $U_{R}$ Profile of $U_{R}$ (12) U-P_R_FollowerNum The number of followers of the Recipient $U_{R}$.
U-P_S_FollowR A binary label indicating $U_{S}$ is following $U_{R}$ (1) or not (0).
U-P_R_FollowS A binary label indicating $U_{R}$ is following $U_{S}$ (1) or not (0).
Data(U-HA) Popularity of $U_{S}$ (4) U-HA_S_LikedRate Avg. # of ‘Likes’ received by latest 50 posts of $U_{S}$.
Public metrics of historical posts of $U_{S}$ and $U_{R}$ Popularity of $U_{R}$(4) U-HA_R_LikedRate Avg. # of ‘Likes’ received by latest 50 posts of $U_{R}$.
U-HA (38) Activity of $U_{S}$ (7) U-HA_S_InteractivePer Percentage of reposts in $U_{S}$’s latest 50 posts.
Activity of $U_{R}$(7) U-HA_R_InteractivePer Percentage of reposts in $U_{R}$’s latest 50 posts.
Interaction (16) U-HA_RS_Mention Number of $U_{R}$’s latest 50 posts that mentioned $U_{S}$.
Data(U-HM) U-HM (157) M features of $U_{S}$ (78) U-HM_S_Grammar1 (Each M feature is averaged over $U_{S}$’s latest 50 posts.)
Text content of historical posts of $U_{S}$ and $U_{R}$ M features of $U_{R}$ (78) U-HM_R_Grammar1 (Each M feature is averaged over $U_{R}$’s latest 50 posts.)
Topic Similarity (1) U-HM_SR_TopicSim Similarity between historical topics of 50 latest posts
of $U_{S}$ and $U_{R}$.
Table 1: Types and sub-types of features used in this study. Features are extracted from post data, Data(M), and user data, Data(U-P), Data(U-HA), and Data(U-HM).
@@ 2.1 Reposting Prediction
A reposting action involves three elements.
  • The content of a post ($M$). So far most works focus on the text content.
  • The sender ($U_{S}$), who created the post.
  • The recipient ($U_{R}$), i.e., the potential reposter, whose action is to be predicted.
In this study, we define reposting prediction as to predict whether a user (recipient) will repost a given post created by another user (sender), considering all of the three elements.
@@
$f(M,U_{S},U_{R})\rightarrow y\in\{0,1\}\,$, (1)
where $y=1$ means the recipient will repost the post received from the sender, and $y=0$ otherwise.
2.2 Related Works
There are many related works on various aspects of reposting. Some works studied posts only, some focused on posts and their senders only, and some considered posts and their recipients only. For example, [36, 63] predicted the popularity of a post; [32, 52] predicted the incremental size of the spreading tree for a given post over time; [57] predicted the time gap between reposts of a given post; [49] predicted the probability of a given post being reposted (by any user), and [48, 26, 11, 58, 55] predicted the next user who is going to repost a given post based on the sequence of previous reposters of the post and the social relations among these reposters.
2.3 Prior Works on Reposting Prediction
Prior works on reposting prediction as defined in Equation (1) include a number of machine learning (ML) models, such as Support Vector Machines (SVM) [53, 34], Decision Trees  [53, 20, 65, 56], Logistic Regression  [59, 42], and factorisation models  [19, 27, 28, 22]. ML models are based on features extracted from raw data of posts and users. It is suggested in [53, 65] that the Decision Tree models often outperform other ML models.
@@ 2.3 Prior Works on Reposting Prediction
Reposting prediction has also been studied from the perspective of recommender systems, which predict and rank the likelihood of candidate posts being reposted by a user [54, 19, 15, 1]. These studies used various prediction models, typically considered all the three elements, and evaluate their model performance at the ranked-list level.
2.4 Post Data vs User Data
Prior works on reposting prediction primarily relied on data of posts (including posts being reposted and the historial posts of users). For example, the ML models typically relied on features extracted from the textual content of post data, and the NN models often used the post data exclusively.
It is known, however, that user-related data, including users’ profile, their following list, and their past behaviour, interest and interactions, can have a significant influence on a user’s status, actions and decisions [34, 30, 21, 53, 61]. For example, [34] suggested that a recipient is more likely to repost a post from a sender if the recipient follows the sender, and the two users have shared interests and prior interactions. Prior works on recommender systems support similar observations  [9, 61, 51].
2.5 Out-of-Distribution Evaluation
Prior works on reposting prediction typically evaluated model performance by splitting the data into training and test sets using random sampling, where both sets contain data from the same set of topics [53, 34, 20, 65, 56, 59, 19, 27, 28, 22, 60, 35]. This setting, often referred to as in-distribution evaluation [3, 33], assumes that the training and test data are drawn from the same underlying distribution.
However, different topics may associate with user groups with distinct feature distributions. None of prior works explicitly evaluated generalisation across different topics in the out-of-distribution setting, that is to train a model using data of some topics (characterised by a set of hashtags) and then test the model using data of other topics, which can evaluate a model’s capacity in terms of predicting reposts associated with new, previously unseen topics.
3 Our Data for Reposting Prediction
@@ 3.1 X Data
We collected data from X using the former Twitter API with the Academic Access. We selected the 14 most popular hashtags in the trending list between July 27th and August 14th in 2022 and retrieved all available posts containing these hashtags during the same period around the world.
We obtained 63,336 posts, including 44,014 reposts (positive instances). These reposts satisfy two conditions: (1) each repost is created within 24 hours since its original post was posted; (2) the original post is also in our data. We set a 24-hour window because, according to the study of the half-life of X posts [41], users are unlikely to see a post more than 24 hours after its creation. Any reposting after 24 hours is likely influenced by external factors, such as TV or other social media platforms.
These posts (including the reposts) were created by 54,299 unique users, which include all senders and recipients of the reposts. For each user, we collected their profile, following list, and up to 50 latest historical posts. Work in [18, 4] suggest that 50 historical posts are sufficient to effectively characterise a user. A total of more than 2.6 million historical posts were collected. All users are anonymised in our data.
3.2 Post Data and Features
A post can have text and multimedia content, such as pictures, audio, and video. Same as in prior works, in this study, we only consider the text content of a post.
  • Data(M): the post data, contain the text content of the posts in our X data.
From the post data, we extract 78 post-related features, M, which are used in prior works for reposting prediction.
These features characterise a post from different perspectives. As shown in Table 1, we divide them into sub-types, including topic  [2, 23, 37, 65, 20, 22], language style [7, 19, 29, 12], readability [29], sentiment [25, 44, 62, 20], emotion [40, 44, 50, 20], hate speech  [38], and hashtag. These features are calculated using publicly available language models and tools (see Appendix). In addition, the hashtag feature is recorded as a categorical label [19, 20].
@@ 3.3 User Data and Features
In this study, the user data, Data(U), contain the following three types of user-related data.
@@
  • Data(U-HA): the metadata (or public metrics) of the latest 50 historical posts of a user, including (1) the type of a post, i.e., original post/repost/quote/reply, (2) the creation time of a post, (3) the number of reposts/quotes/replies/likes a post has received, and (4) any usernames mentioned in a post. These metadata encode useful information of a user’s previous behaviours, activity patterns, and its interactions with other users.
  • Data(U-HM): the textual content of the latest 50 historical posts of a user, which characterise a user’s historical interests and communication style.